OpenStudy (anonymous):

(8x + 8) + (6x - 4) Add

OpenStudy (anonymous):

\[= 14x + 4 = 2(7x + 2)\]

OpenStudy (anonymous):

14x + 4 is what I got

OpenStudy (anonymous):

excellent, simplify and you'll get = 2(7x + 2)

OpenStudy (anonymous):

hey sstarica can u help me as soon as ur done with him...Thanks

OpenStudy (anonymous):

14x +4